OpenStudy (anonymous):

Solve the system y = x + 2 and y = –x – 3 by using graph paper or graphing technology. i got (–2.5 , –0.5) am i right or wrong? please help:)

OpenStudy (anonymous):

y = x + 2 and y = –x – 3 substitute –x – 3 = x + 2 + x +x -3 = 2x + 2 -2 -2 -5 = 2x x=-5/2 =-2.5 back to y = x + 2 y = (-2.5) + 2 = -0.5 or y = –x – 3 y = –(-2.5) – 3 =2.5 - 3= -0.5 You got it